Gud day please help me to solve this:

1.1. Solve the equation -4k2-8k-3=-3-5k2
1.2 factorise 6p2+11pq+3q2
2. Duuring the global leadership convention,it is given that the number of men who are attending the convention is nine hundred and forty less than four times the number of women in attendance.
2.1. from the statement above.formulate a mathematical equation expresing the number of men in terms of women ,given that w represent the number of women and m represent the number of men.
2.2. given that the total number of people who are attending the global leadership convention are two thousand five hundred and sixty,use the equation you formulated in 2.1 to determine the number of men who are attending the convention.

Answer by Boreal(15235)

-4k2-8k-3=-3-5k2
assume this is a quadratic
k^2-8k=0
k(k-8)=0
k=0, 8
--------------------
second is (3p+q)(2p+3q)
-------------------
last is
w=women
4w-940=men
their sum = 2560
w+4w-940=2560
5w=3500
w=700 women
m=2800-940 or 1860, and that adds up to 2560. ANSWER